# Importance Guided Data Augmentation for Neural-Based Code Understanding

### v1

@@ Abstract
Pre-trained code models lead the era of code intelligence. Many models have been designed with impressive performance recently. However, one important problem, data augmentation for code data that automatically helps developers prepare training data lacks study in the field of code learning. In this paper, we introduce a general data augmentation framework, GenCode, to enhance the training of code understanding models. GenCode follows a generation-and-selection paradigm to prepare useful training codes. Specifically, it uses code transformation techniques to generate new code candidates first and then selects important ones as the training data by importance metrics. To evaluate the effectiveness of GenCode with a general importance metric – loss value, we conduct experiments on four code understanding tasks (e.g., code clone detection) and three pre-trained code models (e.g., CodeT5). Compared to the state-of-the-art (SOTA) code augmentation method, MixCode, GenCode produces code models with 2.92% higher accuracy and 4.90% robustness on average.
1 Introduction
@@
Deep learning (DL) for code has been one of the successful DL applications over the years. The methodology of applying deep learning to solve software engineering (SE) tasks, such as software vulnerability detection Zhou et al. (2019), is commonly known as code learning. From the model level, many DL models specifically designed for code learning have been proposed, such as CodeBERT Feng et al. (2020), GraphCodeBERT Guo et al. (2020), and CodeT5 Wang et al. (2021). From the business perspective, multiple related products such as Copilot  [^1] https://github.com/features/copilot and Tabnine  [^2] https://www.tabnine.com have been widely used to help our daily development. As a result, in different downstream learning-based code understanding tasks, e.g., code clone detection  Lu et al. (2021) and program classification Puri et al. (2021), DL already achieved significantly better results than traditional SE techniques.
@@
However, most of the existing research focuses on designing new code representation techniques or code model architectures to further improve the effectiveness of DL for code, one important problem – how to prepare high-quality training data gained limited attention even though recent research demonstrated it highly affects the performance of code models Sun et al. (2022). The challenge behind this is preparing training data for code models is heavy work. Although today developers can easily collect raw code data from the open-source community such as GitHub [^3] https://github.com, data labeling is still expensive, especially for code-related tasks. It needs human effort with professional domain knowledge.
@@
Techniques can be directly used for code augmentation, such as code refactoring methods and text transformation methods. Unfortunately, existing studies Bielik and Vechev (2020); Yu et al. (2022) stated that code refactoring methods have limited advantages in boosting the training of code models. Besides,  Dong et al. (2023b) revealed that almost all the data augmentation methods cannot produce code models with good robustness. Moreover, a recent work, MixCode Dong et al. (2023a), that follows the spirit of Mixup Zhang et al. (2018) to do code augmentation achieved promising results, but it can be only used for classification tasks. Therefore, there are no effective and reliable code augmentation methods that can be generalized to different tasks.
@@
To bridge this gap, inspired by the great success of the typical two-stage Automated Data Augmentation (AutoDA) framework in the field of computer vision Cubuk et al. (2020); Yang et al. (2023b), we propose GenCode, the first generation-and-selection-based data augmentation framework for code understanding. Specifically, at each training epoch, GenCode first generates training candidates by using different code transformation methods. Here, we consider different types of code transformation methods such as semantic-preserving code refactoring methods and syntax-breaking text transformation methods. After that, GenCode selects valuable training data from the candidates with the guidance of importance metrics. The important metric used in GenCode is simply defined as the loss value of code data produced by the code model. Finally, the selected samples are used for the model training. We conducted comprehensive experiments to study how GenCode can improve the accuracy and robustness of code models. The experimental results demonstrated that, compared to the SOTA method, MixCode, GenCode can produce code models with a 2.92% higher accuracy and 4.90% better robustness on average.
@@
  • We propose the first generation-and-selection-based framework, GenCode, which automatically enhances the training data and identifies the most important data from the augmented set for boosting the code understanding.
@@
2 Background and Related Work
2.1 Deep Learning for Code
@@
Since source code can be seen as natural language or structural graph Allamanis et al. (2018), researchers employ the techniques in Natural Language Processing (NLP) and Graph Neural Networks (GNNs) to learn information from code and solve downstream tasks, such as automated bug detection Dinella et al. (2020), automated program classification Puri et al. (2021), and code clone detection Wang et al. (2020). The key problem in deep learning for code is how to represent raw code to machine-readable inputs. Researchers designed different methods for code representation that can be roughly divided into two groups, sequence-based representation and graph-based representation.
Sequence-based representation follows the same procedure of text data processing in NLP to convert raw code data into a sequence of code tokens. For example, given a code snippet, “def func(x, y)”, sequence-based representation transforms it to a token list as “[’def’, ’func’, ’(’, ’x’, ’y’, ’)’]”. After that, each token in the token list will be mapped to a pre-defined index value, e.g., [1, 25, 34, 5100, 640, 33]. Finally, these index values will be fed into the code-embedding models for the learning.
Graph-based representation learns the structural information of code. Here, the structural information can be abstract syntax tree (AST), data flow information, and other types of flow information Dinella et al. (2020); Yasunaga and Liang (2020). For example, GraphCodeBERT combines the sequence-based representation and data flow information of code as a whole to train code models with better performance.
@@
In this work, we focus on both sequence-based and graph-based representation and study the powerful pre-trained code models CodeBERT, GraphCodeBERT, and CodeT5 that achieved impressive results on code understanding tasks.
2.2 Data Augmentation for Code
@@
Data augmentation is a well-studied technique in the machine learning field to increase the size and diversity of training data by only modifying the existing labeled data, thus, without further human effort. For instance, in the NLP field, researchers tried many ways to augment the text data, e.g., replacing a token with a synonym Feng et al. (2021).
@@
Recently, data augmentation gained attention from software engineering researchers. Various data augmentation methods designed for source code have been proposed Gao et al. (2020); Yefet et al. (2020); Allamanis et al. (2021); Pour et al. (2021); Bui et al. (2021); Wang et al. (2022); Park et al. (2023); Yang et al. (2023a), with the most general framework involving the use of code refactoring, e.g., local variable renaming, duplication, and dead store, to create more code data without breaking the semantics of the original codes. Concretely, code refactoring rewrites the syntactic structure of source code while keeping the semantic information Lacerda et al. (2020).
@@
Existing works Bielik and Vechev (2020); Yu et al. (2022); de Paula Rodrigues et al. (2023); Dong et al. (2023b, 2024) conducted empirical studies and found code refactoring methods have limited advantages in boosting the training of code models. For instance, a recent empirical study Dong et al. (2024) on code understanding revealed that the utilization of code refactoring methods only achieved success rates of 66.67% and 41.67% in improving accuracy and robustness, respectively, compared to code model training without data augmentation. Besides, the SOTA code data augmentation method, MixCode which linearly interpolates pairs of source codes, has been proven more effective than the current general framework code refactoring. However, as a variant of Mixup specifically designed for code data, MixCode is limited to applications in code classification tasks, as it necessitates the label with one-hot encoding.
@@ 3 Definition
Definition 3.1 (Program Code [^4] For simplicity, we use code in the paper. ).
A code is defined as a 5-tuple $\mathcal{P}=\left(content,se,sy,Ip,Op\right)$: $content$ contains all the text in the code $\mathcal{P}$; $se$ and $sy$ are the semantic and syntax of the code $\mathcal{P}$; and $Ip$ and $Op$ are the input and output of the code $\mathcal{P}$.
Definition 3.2 (Code Model).
@@
Code models can be any machine learning model such as a simple regression model or a pre-trained model, e.g., CodeT5 Wang et al. (2021).
Definition 3.3 (Code Augmentation).
Given a code $P$, a code transformation method $\mathcal{T}:\mathcal{X}\rightarrow\mathcal{X^{\prime}}$, code augmentation is to transform $\mathcal{P}$ to $\mathcal{P^{\prime}}$ that $\mathcal{P^{\prime}}.content=\mathcal{T}\left(\mathcal{P}.content\right)$.
@@
Definition 3.4 (Semantic-preserving/breaking Code Data Augmentation).
@@
Definition 3.5 (Syntax-preserving/breaking Code Augmentation).
@@
4 Assumption and Preliminary Study
GenCode selects the most valuable data for the code model training. This approach is grounded on the assumption that the data that code models have less confidence in is more useful for model training. To check whether this assumption is valid or not, we conducted a preliminary study to explore if there is a correlation between the accuracy of the trained code model and the confidence of the model on the training data. Specifically, the study is conducted on the Google Code Jam (GCJ) Yang et al. (2022) dataset and CodeBERT pre-trained code model. The detailed steps are as follows:
Step 1. We initialize the code model and use code augmentation methods to generate additional data as the training candidate set. The code augmentation methods will be introduced in Section 5.2.
Step 2. We randomly select a number of data from the training candidate set. Here the selected number is the same as the size of the original training data. We repeat this selection 100 times and collect 100 groups of training data. After that, for each group of training data, we compute and record their average loss values.
@@ 4 Assumption and Preliminary Study
Step 4. We use the Pearson Correlation Coefficient (PCC) Schober et al. (2018) metric to check the correlation between the loss values and the accuracies.
Table 1: Correlation between loss values and code model accuracy.
@@
Table 1 summarizes the results of our preliminary study. Here, we considered three types of code models, the initialed model, the model that has been trained for a few epochs but not converged, and the well-trained code model, to check if our assumption can stand in different scenarios. The results indicate that there is a positive correlation between the loss values of data and the accuracy of code models trained by using these data. From the initialized model to the well-trained model, this correlation becomes weaker but still exists. Therefore, we can propose an effective and reliable data augmentation framework to boost the training of code models based on this phenomenon.
@@ 5.1 Overview
GenCode follows the generation-and-selection workflow and utilizes different types of code transformation methods to generate diverse code data first, then selects code data that are most useful for training as the training data. Algorithm 1 shows how GenCode works. In each training epoch, GenCode automatically generates more code data (Line 7 to Line 9, Definition 3.3) and then selects the most important ones (Line 10 to Line 13) to train the code model (Line 14). Here, the data importance is measured by the loss values produced by the code model which is the most common indicator to check how the model performs on the data. Overall, GenCode contains two main steps, utilizing code data augmentation methods to compose the search space and importance-guided data selection.
@@
5.2 Search Space
Even though the code has strict syntax constraints, recent work Dong et al. (2023b) found that these data augmentation methods that slightly break the syntax of the source code e.g., Random Swap, can still be useful for code model training. To enhance the diversity of source code within the feature space, therefore, GenCode considers two types of code augmentation methods, semantic/syntax-preserving methods, and syntax-breaking methods (see Definition 3.4 and Definition 3.5).
GenCode adapts code refactoring methods as semantic/syntax-preserving methods. Code refactoring methods follow the constraints of programming languages to edit the code while keeping the main functionality of the code. On the other hand, text augmentation methods are used as syntax-breaking methods in GenCode which are widely used in the NLP field. For example, text augmentation methods can randomly delete some tokens in the code. However, even though the syntax has been broken, the code model can still successfully predict those codes since it reads the code as a sequence. In total, as shown in Table 2, GenCode currently supports 18 code refactoring methods and 5 text augmentation methods. Besides, Figure 1 provides an example of the training candidate that is to generate the search space. In the syntax-preserving, we present a code refactoring method named API Renaming, which involves modifying the function name from ”main()” to ”even()”. In the syntax-breaking, we demonstrate another instance of API Renaming, where two different statements within a program are randomly selected and their positions swapped. As depicted in this example, the third and fourth lines have been interchanged. Appendix B detailed introduces the specifics of these atomic data augmentation operators.
@@ 5.2 Search Space
Synonym replacement Random insertion Random swap Random deletion Back-translation
@@
Model DA method Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$
@@
5.3 Importance-Guided Data Selection
@@
After constructing the training candidate (as introduced in Section 5.2), the next step is to choose the most valuable code data that can benefit the code model training. Based on our Finding drawn from the preliminary study in Section 4, GenCode chooses the simplest way to measure the importance of code data to the model – the loss value. That means, GenCode takes all the transformed code as input to the code model and records all the loss values first, then ranks all the code in descending order based on the loss values and selects the first $K$ code as the training data. Following the common setting in the data augmentation problem Dong et al. (2023a), $K$ is set as the number of original training data. In GenCode, we choose loss value as the indicator since it can be generalized to any code model and task.
6 Evaluation
@@
We evaluate both the accuracy (Section 6.2) and robustness (Section 6.3) of trained code models. Besides, we check how does importance score affects the effectiveness of GenCode (Section 6.4).
6.1 Datasets and Models
@@
Datasets and Models We evaluate GenCode on four code understanding tasks, bug detection Hu et al. (2019), authorship attribution Yang et al. (2022), code clone detection Yang et al. (2022) (Note that the SOTA method, MixCode, is not suitable for this task since its input is code pairs.), and problem classification Puri et al. (2021). For each code task, we fine-tune three types of models based on pre-trained code models, CodeBERT, GraphCodebERT, and CodeT5. For the model training, we follow existing works Hu et al. (2019); Wang et al. (2021); Puri et al. (2021); Yang et al. (2022); Dong et al. (2023a), to set all configurations. Specifically, we set the training epoch as 50 for all tasks and models. Regarding the optimizer, we utilize Adam Kingma and Ba (2014) with a learning rate of $10^{-3}$ for all the models. Besides, we use early stopping with a patience of 20 to mitigate overfitting. To reduce the influence of randomness, we train each model five times and report the average results along with the standard deviation. For MixCode, we follow the recommendation of the original paper and set the Mixup ratio as $\alpha=0.1$. For the setting of adversarial attacks, we directly use the default configurations provided by the original papers  Yang et al. (2022) in our work. More details are explained in Appendix A and C.
@@
Evaluation Metrics. In the evaluation, we assess the performance of trained models from two perspectives, namely, clean accuracy and model robustness. For the robustness evaluation, we check the attack success rate (ASR) metric of two well-established adversarial attack methods, namely, ALERT Yang et al. (2022) and MHM Zhang et al. (2020) on code models. We do not consider the natural robustness since the existing natural robustness datasets can not fully support our studied tasks, i.e., no natural robustness datasets support authorship attribution, bug detection, and code clone detection.
@@
Baseline Methods. We consider model training without data augmentation (No Aug), Refactor, a family of text-oriented data augmentation methods, and MixCode as baseline methods. Refactor is the widely used general data augmentation framework that has been specifically designed for code data, while MixCode is the more recent code data augmentation method with high performance in enhancing code models Dong et al. (2023a). These baselines allow us to determine if there are real improvements in the accuracy and robustness of the code model when using GenCode.
6.2 Accuracy Evaluation
@@
Overall accuracy. Column Accuracy in Table 3 summarizes the final test accuracy of code models trained by using different code data augmentation methods including GenCode. Surprisingly, from the results, we can see that most of the existing code augmentation methods except RS and MixCode negatively affect the performance of code models. This finding is similar to the conclusion from recent code augmentation studies Bielik and Vechev (2020); Yu et al. (2022); Dong et al. (2023b) which states code augmentation has limited advantage in boosting the training of code models and also indicates that data augmentation for code is a challenging problem.
@@
Then, compared to the existing data augmentation methods, the results demonstrate that GenCode consistently enhances the training and achieves the best results across all the considered datasets and models. Concretely, on average, GenCode generates models with 4.52%, 3.05%, 2.56%, and 2.11% higher accuracy than No Aug in GCJ, Refactory, Java250, and BigCloneBench, respectively. Besides, compared to the SOTA code augmentation method MixCodde, GenCode can achieve 4.1%, 2.34%, and 2.33% better results in GCJ, Refactory, and Java250 datasets, respectively. In conclusion, GenCode is effective in training accurate code understanding models.
@@
Case study 1: Does the training epoch affect the accuracy of GenCode? Figure 2 depicts the average test accuracy at each training epoch of CodeBERT using different data augmentation methods. The results demonstrated that GenCode consistently outperforms other baseline methods during the whole training process. We can conclude that the choice of the number of training epochs will not affect the advantages of our method over other methods.
@@
Table 4: A case study of mean values of maximum probabilities. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
@@
Model DA method Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$

### v2

# GenCode: A Generic Data Augmentation Framework for Boosting Deep Learning-Based Code Understanding

@@ Abstract
Pre-trained code models lead the era of code intelligence with multiple models have been designed with impressive performance. However, one important problem, data augmentation for code data that automatically helps developers prepare training data lacks study in this field. In this paper, we introduce a generic data augmentation framework, GenCode, to enhance the training of code understanding models. Simply speaking, GenCode follows a generation-and-selection paradigm to prepare useful training code data. Specifically, it employs code transformation techniques to generate new code candidates first and then selects important ones as the training data by importance metrics. To evaluate the effectiveness of GenCode, we conduct experiments on four code understanding tasks (e.g., code clone detection) and three pre-trained code models (e.g., CodeT5). Compared to the state-of-the-art (SOTA) code augmentation method, MixCode, GenCode produces code models with 2.92% higher accuracy and 4.90% robustness on average.
Introduction
@@
Deep learning (DL) for code has been one of the successful DL applications over the years. The methodology of applying deep learning to solve software engineering (SE) tasks, such as software vulnerability detection (Zhou et al. 2019), is commonly known as code learning. From the model level, different DL models specifically designed for code learning have been proposed, such as CodeBERT (Feng et al. 2020), GraphCodeBERT (Guo et al. 2020), and CodeT5 (Wang et al. 2021). From the commercial perspective, multiple related products such as Copilot  [^1] https://github.com/features/copilot and Tabnine  [^2] https://www.tabnine.com have been widely used to help our daily development. As a result, in different downstream learning-based code understanding tasks, e.g., code clone detection  (Lu et al. 2021) and program classification (Puri et al. 2021), DL already achieved significantly better results than traditional SE techniques.
@@
However, most of the existing research focuses on designing new code representation techniques or code model architectures to further improve the effectiveness of DL for code, one important problem – how to prepare high-quality training data gained limited attention even though recent research demonstrated it highly affects the performance of code models (Sun et al. 2022). The challenge behind this is preparing training data for code models is heavy work. Although today developers can easily collect raw code data from the open-source community such as GitHub  [^3] https://github.com, data labeling is still expensive, especially for code-related tasks. It needs huge human effort with professional domain knowledge.
@@
To tackle this problem, data augmentation is a famous technique to increase the diversity and quality of training data based only on existing labeled data. In the code learning field techniques such as code refactoring methods and text-oriented transformation methods can be directly used for code augmentation. Unfortunately, existing studies (Bielik and Vechev 2020; Yu, Wang, and Wang 2022) stated that code refactoring methods have limited advantages in the accuracy improvement of code models. Besides,  (Dong et al. 2023b) revealed that code models trained using existing data augmentation methods still suffer from the low robustness issue. Moreover, a recent work, MixCode (Dong et al. 2023a), that follows the spirit of Mixup (Zhang et al. 2018) to do code augmentation achieved promising results, but it can be only used for classification tasks. Therefore, there are currently no effective and reliable code augmentation methods that can be generalized across different tasks.
@@
To bridge this gap, inspired by the great success of the typical two-stage Automated Data Augmentation (AutoDA) framework in the field of computer vision (Cubuk et al. 2020; Yang et al. 2023b), we propose GenCode, the first generation-and-selection-based data augmentation framework for code understanding. Specifically, at each training epoch, GenCode first generates training candidates by using different code transformation methods. Here, we consider different types of methods such as semantic-preserving code refactoring methods and syntax-breaking text transformation methods. After that, GenCode selects valuable training data from the candidates with the guidance of importance metrics. The important metric used in GenCode is simply defined as the loss value of code data produced by the code model. Finally, the selected samples are used for the model training. We conducted comprehensive experiments to study how GenCode can improve the accuracy and robustness of code models. The experimental results demonstrated that, compared to the SOTA method, MixCode, GenCode can produce code models with a 2.92% higher accuracy and 4.90% better robustness on average.
@@
  • We propose the first generation-and-selection-based framework, GenCode, which automatically enriches the training data for boosting the code understanding models.
@@
Background and Related Work
Code Representation
@@
Code representation is a crucial technique that converts source code into a format interpretable by Deep Neural Networks (DNNs), facilitating effective learning of source code features. Since source code can be seen as sequences of tokens or structural graph (Allamanis et al. 2018), researchers employ the techniques in Natural Language Processing (NLP) and Graph Neural Networks (GNNs) to learn information from code and solve downstream tasks, such as automated program classification (Puri et al. 2021) and code clone detection (Wang et al. 2020). Existing methods for code representation can be roughly divided into two groups, sequence-based representation and graph-based representation. Sequence-based representation follows the same procedure of text data processing in NLP to convert raw code data into a sequence of code tokens (Alon et al. 2019). Graph-based representation learns the structural information of code. Here, the structural information can be abstract syntax tree (AST), data flow information, and other types of flow information (Dinella et al. 2020; Yasunaga and Liang 2020).
@@
Unlike code representation, our work focuses on code data preprocessing and studies the powerful pre-trained code models CodeBERT, GraphCodeBERT, and CodeT5 that achieved impressive results on code understanding tasks.
Data Augmentation for Code Understanding
@@
Data augmentation is a well-studied technique in the machine learning field that increases the size and diversity of training data by only modifying the existing labeled data, thus, without further human effort. For instance, in the NLP field, researchers tried many ways to augment the text data, e.g., replacing a token with a synonym (Feng et al. 2021).
@@
Recently, data augmentation has gained attention from software engineering researchers. Various data augmentation methods designed for source code have been proposed (Gao et al. 2020; Yefet, Alon, and Yahav 2020; Allamanis, Jackson-Flux, and Brockschmidt 2021; Pour et al. 2021; Bui, Yu, and Jiang 2021; Wang et al. 2022; Park, Kim, and Han 2023; Yang et al. 2023a), with the most general framework involving the use of code refactoring, e.g., local variable renaming and dead store, to create more code data without breaking the semantics of the original codes. Concretely, code refactoring rewrites the syntactic structure of source code while keeping the semantic information (Lacerda et al. 2020).
@@
Existing works (Bielik and Vechev 2020; Yu, Wang, and Wang 2022; de Paula Rodrigues, Braga, and Dahab 2023; Dong et al. 2023b, 2024) conducted empirical studies and found code refactoring methods have limited advantages in boosting the training of code models. For instance, a recent empirical study (Dong et al. 2024) on code understanding revealed that the utilization of code refactoring methods only achieved success rates of 66.67% and 41.67% in improving accuracy and robustness, respectively, compared to code model training without data augmentation. Besides, the SOTA code data augmentation method, MixCode, which linearly interpolates pairs of source codes, has been proven more effective than the current general framework code refactoring. However, as a variant of Mixup specifically designed for code data, MixCode is limited to applications in code classification tasks, as it necessitates the label with one-hot encoding.
@@
Definition 1.
A program code  [^4] For simplicity, we use code in the paper. is defined as a 5-tuple $\mathcal{P}=\left(content,se,sy,Ip,Op\right)$: $content$ contains all the text in the code $\mathcal{P}$; $se$ and $sy$ are the semantic and syntax of the code $\mathcal{P}$; and $Ip$ and $Op$ are the input and output of the code $\mathcal{P}$.
Definition 2.
@@
Code models can be any machine learning model such as a simple regression model or a pre-trained model, e.g., CodeT5 (Wang et al. 2021).
Definition 3.
Given a code $P$, a code transformation method $\mathcal{T}:\mathcal{X}\rightarrow\mathcal{X^{\prime}}$, code augmentation is to transform $\mathcal{P}$ to $\mathcal{P^{\prime}}$ that $\mathcal{P^{\prime}}.content=\mathcal{T}\left(\mathcal{P}.content\right)\&\mathcal{P^{\prime}}.content\neq\mathcal{P}.content$.
@@
Definition 4.
@@
Definition 5.
@@
Assumption and Preliminary Study
GenCode selects the most valuable data for the code model training. This approach is grounded on the assumption that the data that code models have less confidence in is more useful for model training. To check whether this assumption is valid or not, we conducted a preliminary study to explore if there is a correlation between the accuracy of the trained code model and the confidence of the model on the training data. Specifically, the study is conducted on the Google Code Jam (GCJ) (Yang et al. 2022) dataset and CodeBERT pre-trained code model. The detailed steps are as follows:
Step 1. We initialize the code model and use code augmentation methods to generate additional data as the training candidate set. The code augmentation methods will be introduced in a later section.
Step 2. We randomly select a number of data from the training candidate set. Here the selected number is the same as the size of the original training data. We repeat this selection 100 times and collect 100 groups of training data. After that, we compute and record the average loss values for each group of training data.
@@
Step 4. We use the Pearson Correlation Coefficient (PCC) (Schober, Boer, and Schwarte 2018) metric to check the correlation between the loss values and the accuracies.
Table 1: Correlation between loss values and code model accuracy. The P-value less than 0.01 is highlighted with a blue background.
@@
Table 1 summarizes the results of our preliminary study. Here, we considered three types of code models, the initialed model, the model that has been trained for a few epochs but not converged, and the well-trained code model, to check if our assumption can stand in different scenarios. The results indicate that there is a positive correlation between the loss values of data and the accuracy of code models trained by using these data. From the initialized model to the well-trained model, this correlation becomes weaker but still exists. We conjecture that when the model is in the early stage, most of the data that have high loss values are useful for the model training. However, when training proceeds, the data that still have high loss values are the model itself cannot learn. Thus, there will be a correlation drop from the early epoch to the later ones. Based on this phenomenon, we can propose an effective and reliable data augmentation framework to boost the training of code models based on this phenomenon.
@@
GenCode follows the generation-and-selection workflow and utilizes different types of code transformation methods to generate diverse code data first, then selects code data that are most useful for training as the training data. Algorithm 1 shows how GenCode works. In each training epoch, GenCode automatically generates more code data (Line 7 to Line 9, Definition 3) and then selects the most important ones (Line 10 to Line 13) to train the code model (Line 14). Here, the data importance is measured by the loss values produced by the code model which is the most common indicator to check how the model performs on the data. Overall, GenCode contains two main steps, utilizing code data augmentation methods to compose the search space and importance-guided data selection.
@@
Search Space
Even though the code has strict syntax constraints, recent work (Dong et al. 2023b) found that these data augmentation methods that slightly break the syntax of the source code e.g., Random Swap, can still be useful for code model training. To enhance the diversity of source code within the feature space, therefore, GenCode considers two types of code augmentation methods, semantic/syntax-preserving methods, and syntax-breaking methods (see Definition 4 and Definition 5).
GenCode adapts code refactoring methods as semantic/syntax-preserving methods. Code refactoring methods follow the constraints of programming languages to edit the code while keeping the main functionality of the code. On the other hand, text augmentation methods are used as syntax-breaking methods in GenCode which are widely used in the NLP field. For example, text augmentation methods can randomly delete some tokens in the code. However, even though the syntax has been broken, the code model can still successfully predict those codes since it reads the code as a sequence. In total, as shown in Table 2, GenCode currently supports 18 code refactoring methods and 5 text augmentation methods. Besides, Figure 1 provides an example of the training candidate that is to generate the search space. In the syntax-preserving, we present a code refactoring method named API Renaming, which involves modifying the function name from ”main()” to ”even()”. In the syntax-breaking, we demonstrate another instance of API Renaming, where two different statements within a program are randomly selected and their positions swapped. As depicted in this example, the third and fourth lines have been interchanged.
@@
Synonym replacement (SR) Random insertion (RI) Random swap (RS Random deletion (RD) Back-translation (BT)
@@
Model DA method Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$
@@
Importance-Guided Data Selection
@@
After constructing the training candidate, we then choose the most valuable code data that can benefit the code model training. Based on our Finding drawn from the preliminary study, GenCode chooses the simplest way to measure the importance of code data to the model – the loss value. That means, GenCode takes all the transformed code as input to the code model and records all the loss values first, then ranks all the code in descending order based on the loss values and selects the first $K$ code as the training data. Following the common setting in the data augmentation problem (Dong et al. 2023a), $K$ is set as the number of original training data. GenCode chooses loss value as the indicator since it can be generalized to any code model and task.
Evaluation
@@
We evaluate both the accuracy and robustness of trained code models. Besides, we check how does importance score affects the effectiveness of GenCode.
Datasets and Models
@@
Datasets and Models We evaluate GenCode on four code understanding tasks, bug detection (Hu et al. 2019), authorship attribution (Yang et al. 2022), code clone detection (Yang et al. 2022) (Note that, MixCode is not suitable for this task since its input is code pairs.), and problem classification (Puri et al. 2021). For each code task, we fine-tune three types of models based on pre-trained code models, CodeBERT, GraphCodebERT, and CodeT5. For the model training, we follow existing works (Hu et al. 2019; Wang et al. 2021; Puri et al. 2021; Yang et al. 2022; Dong et al. 2023a), to set all configurations. Specifically, we set the training epoch as 50 for all tasks and models. Regarding the optimizer, we utilize Adam (Kingma and Ba 2014) with a learning rate of $10^{-3}$ for all the models. Besides, we use early stopping with a patience of 20 to mitigate overfitting. To reduce the influence of randomness, we train each model five times and report the average results along with the standard deviation. For MixCode, we follow the recommendation of the original paper and set the Mixup ratio as $\alpha=0.1$. For the setting of adversarial attacks, we directly use the default configurations provided by the original papers  (Yang et al. 2022).
@@
Evaluation Metrics. In the evaluation, we assess the performance of trained models from two perspectives, namely, clean accuracy and model robustness. For the robustness evaluation, we check the attack success rate (ASR) metric of two well-established adversarial attack methods, namely, ALERT (Yang et al. 2022), which uses variable renaming while considering natural semantics, and MHM (Zhang et al. 2020), which also randomly replaces variable names but does not account for natural semantics. We do not consider the natural robustness since the existing natural robustness datasets can not fully support our studied tasks, i.e., no natural robustness datasets support authorship attribution, bug detection, and code clone detection.
@@
Baseline Methods. We consider model training without data augmentation (No Aug), Refactor (Yu, Wang, and Wang 2022), a family of text-oriented data augmentation methods (Dong et al. 2023b), and MixCode as baseline methods. Refactor, which keeps code syntax rules, is a widely used data augmentation framework specifically designed for code data. In contrast, text-oriented code data augmentation methods, which slightly break syntax rules, have recently proven more effective than Refactor. Additionally, MixCode is a more recent and high-performing code data augmentation method that significantly enhances code models.
Accuracy Evaluation
@@
Overall accuracy. Column Accuracy in Table 3 summarizes the final test accuracy of code models trained by using different code data augmentation methods including GenCode. Surprisingly, from the results, we can see that most of the existing code augmentation methods except RS and MixCode negatively affect the performance of code models. This finding is similar to the conclusion from recent code augmentation studies (Bielik and Vechev 2020; Yu, Wang, and Wang 2022; Dong et al. 2023b) which states code augmentation has limited advantage in boosting the training of code models and also indicates that data augmentation for code is a challenging problem.
@@
Then, compared to the existing data augmentation methods, the results demonstrate that GenCode consistently enhances the training and achieves the best results across all the considered datasets and models. Concretely, on average, GenCode generates models with 4.52%, 3.05%, 2.56%, and 2.11% higher accuracy than No Aug in GCJ, Refactory, Java250, and BigCloneBench, respectively. Besides, compared to MixCode, GenCode can achieve 4.1%, 2.34%, and 2.33% better results in GCJ, Refactory, and Java250 datasets, respectively.
Statistical analysis. We conducted statistical tests, including the T-Test and Wilcoxon signed-rank test, across all pre-trained models. The results show that GenCode significantly (with p-value $<$ 0.05) outperforms other methods including MixCode. In conclusion, GenCode is effective in training accurate code models.
@@
Convergence speed analysis. Figure 2 depicts the average test accuracy at each training epoch of CodeBERT using different data augmentation methods. The results demonstrated that GenCode consistently outperforms other baseline methods during the whole training process. We can conclude that the choice of the number of training epochs will not affect the advantages of our method over other methods.
@@
Considering Figure 3, we can see that the embedding produced by GenCode is clearly distributed compared to other methods. Besides, we computed the average distance between the data points from different classes, i.e., the average distance between red points and blue points in Figure 3. The results are shown in the following Table. The results confirm that GenCode can better distinguish the two classes. Interestingly, we found that MixCode has the smallest distance result. That means even though MixCode can train code models with good performance, the code models have low confidence in their predictions, thus, the data are near the decision boundaries. This phenomenon is similar to the findings in (Verma et al. 2019), where the representations obtained by interpolating hidden states have two desirable properties: the class-representations are flattened into a minimal amount of directions of variation, and all points in between these flat representations, most unobserved during training and off the data manifold, are assigned low-confidence predictions.
Table 4: Average distance between the data points from different classes. The best is highlighted with a blue background, and the worst is highlighted with a gray background.
DA method Refactor SR RI RS
Average distance 6.02 8.05 7.02 8.06
DA method RD BT MixCode GenCode
Average distance 7.03 6.00 4.05 8.78
Table 5: Mean values of maximum probabilities. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
@@
Model DA method Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$